OpenStudy (anonymous):

What is the slope of the line that passes through (0, 0) and (4, 1)?

OpenStudy (anonymous):

slope=y2-y2/x2-x1 4-0/1-0=4

OpenStudy (anonymous):

First, use this tool, average slope: \[m=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Break your coordinates up, using one set for the 1's and one set for the 2's. Then put your slope back in for y = mx + b But this time, now that you know the slope, put in either of your coordinates in for x & y, replace/substitute. Then solve for b. You'll have both m & b at that point, so just rewrite it as y=(something)x + (something)